OpenStudy (anonymous):

@tHe_FiZiCx99 @Rina.r What are the next three terms in the pattern below? 4, 8, 12, 16, … A. 64, 256, 1024 B. 28, 44, 72 C. 32, 64, 128 D. 20, 24, 28

OpenStudy (anonymous):

@Rina.r

OpenStudy (the_fizicx99):

This is function notation huh?

OpenStudy (anonymous):

no patterns

OpenStudy (the_fizicx99):

D

OpenStudy (alekos):

Its D

OpenStudy (anonymous):

Which equation describes the rule in the pattern below? x 0 3 6 9 12 y 0 12 24 36 48 A. y = x + 3 B. y = 4x C. D. y = x + 12

OpenStudy (anonymous):

D. Adding four each time. Arithmetic sequence.

OpenStudy (the_fizicx99):

B.

OpenStudy (anonymous):

What is the missing term, a, in the pattern below? 28, 14, a, 3.5, 1.75, … A. 0 B. 10 C. 7.5 D. 7

OpenStudy (the_fizicx99):

It's a multiple of 2. 1.75 * 2 = 3.5

OpenStudy (anonymous):

its not on there is it ABCD

OpenStudy (anonymous):

@RIN

OpenStudy (anonymous):

@Rina.r

OpenStudy (the_fizicx99):

It's 7 ...

OpenStudy (the_fizicx99):

3.5*2 = 7 U_U

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

yes u do lol

OpenStudy (anonymous):

What is the missing term, x, in the pattern below? 1, 8, x, 64, 125, … A. 27 B. 15 C. 56 D. 32

OpenStudy (anonymous):

What is the missing term, x, in the pattern below? 1, 8, x, 64, 125, … A. 27 B. 15 C. 56 D. 32

OpenStudy (the_fizicx99):

8 + 56 = 64 however I'm not too sure its C. My guess is C tho

OpenStudy (anonymous):

i just turned it in

OpenStudy (anonymous):

yay i got 80%

OpenStudy (anonymous):

its was not u

OpenStudy (anonymous):

it was me

OpenStudy (anonymous):

ididnt want to put save image so i just guessed on it

OpenStudy (anonymous):

and the 4 i got right from u math master

OpenStudy (the_fizicx99):

Mine were right or wrong?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

i just guessed on 1

OpenStudy (anonymous):

i will so u can get more medals